# Multiple Hypothesis Testing To Estimate The Number Of Communities in Stochastic Block Models

### v1

@@ Abstract
Clustering of single-cell RNA sequencing (scRNA-seq) datasets can give key insights into the biological functions of cells. Therefore, it is not surprising that network-based community detection methods (one of the better clustering methods) are increasingly being used for the clustering of scRNA-seq datasets. The main challenge in implementing network-based community detection methods for scRNA-seq datasets is that these methods apriori require the true number of communities or blocks for estimating the community memberships. Although there are existing methods for estimating the number of communities, they are not suitable for noisy scRNA-seq datasets. Moreover, we require an appropriate method for extracting suitable networks from scRNA-seq datasets. For addressing these issues, we present a two-fold solution: i) a simple likelihood-based approach for extracting stochastic block models (SBMs) out of scRNA-seq datasets, ii) a new sequential multiple testing (SMT) method for estimating the number of communities in SBMs. We study the theoretical properties of SMT and establish its consistency under moderate sparsity conditions. In addition, we compare the numerical performance of the SMT with several existing methods. We also show that our approach performs competitively well against existing methods for estimating the number of communities on benchmark scRNA-seq datasets. Finally, we use our approach for estimating subgroups of a human retina bipolar single cell dataset.
@@ 1.1 Motivation
Recently, several authors have proposed multiple approaches to tackle the problem of clustering of noisy scRNA-seq datasets. Broadly these ideas can be organized into the following steps: i) filtering low-quality cells and genes, ii) log-normalizing the read counts, iii) reducing the number of features or performing dimension reduction, iv) estimating the number of communities, and v) clustering the transformed data; see Heumos et al. (2023) [2] for a thorough discussion. The first step of filtering low-quality genes and cells are required for reducing the noise in scRNA-seq datasets whereas log-normalizing in the second step is required for boosting the signal strength while facilitating an even comparison of cells. Unfortunately, the first two steps cannot completely reduce the elevated noise level from scRNA-seq datasets and further steps are warranted. The existing methodology deals with the remaining noisiness of scRNA-seq datasets either through dimension reduction methods or by estimating the number of blocks. However, the existing approaches require the fine-tuning of hyperparameters, see Shekhar et al. (2016) [3], Kiselev et al. (2017) [4]. For instance, Louvain and Leiden algorithms implemented in Seurat (Satija et al. (2015) [5]), require the number of nearest neighbors beforehand to construct a weighted network for which they estimate the number of communities by modularity maximization approach (Newman and Girvan (2004) [6]). As a consequence, Louvain algorithm tends to result in disjointed clusters see Traag et al. (2019) [7] which is not desirable.
We use a likelihood-based method for constructing a network (with block structures) out of scRNA-seq datasets. Additionally, we propose a sequential multiple testing (SMT) method that is more suited for estimating the number of blocks or communities. We discover that SMT is less sensitive to the noisiness compared to the other existing methods. In particular, we see that the numerical performance of the existing approaches are more sensitive to the ratio of the probability of forming edges between communities to the probability of forming edges within communities, the out-in ratio which is related to the noisiness of the network. Summing up, it is desirable to have a network-based clustering procedure that is suitable for handling the inherent noisiness of extracted networks out of scRNA-seq datasets while requiring minimal or no fine-tuning of free hyperparameters. Our proposed approach is a step in that direction.
@@ 1.2 Stochastic Block Models
The stochastic block model (SBM) proposed by Holland et al. (1983) [8] is one of the popular models for network data with the block or community structure. SBM simplifies the network model by assuming that the edge probability between any pair of nodes is solely determined by their respective community memberships. Although SBM has found favor in many application areas, it is still somewhat not adequate in the other application domains. Recently many extensions and variants of SBM were proposed. For instance, Airoldi et al. (2008) [9] proposed the mixed membership stochastic block model allowing nodes to belong to multiple networks. Karrer and Newman (2011) [10] proposed degree corrected stochastic block models (DCSBMs) relaxing the assumption of homogeneity of nodes. Currently, there are many existing methods for estimating the community structure of SBMs including modularity maximization (Newman and Girvan (2004) [6]), Louvain modularity algorithms (Blondel et al. (2008) [11]), likelihood-based approaches (Bickel and Chen (2009) [12], Zhao et al. (2012) [13], Choi et al. (2012) [14], Amini et al. (2013) [15]), spectral clustering methods (Rohe et al. (2011) [16], Lei and Ronaldo (2015) [17], Joseph and Yu (2016) [18], Sengupta and Chen (2015) [19]) amongst others, see Zhao et al. (2017) [20] for a review. Several existing methods such as Newman and Girvan (2004) [6], Bickel and Chen (2009) [12], Zhao et al. (2012) [13], Qin and Rohe (2013) [21], Amini et al. (2013) [15] have also shown to be consistent for both sparse and dense DCSBMs. However, invariably all of the existing community detection methods for SBMs require the true number of communities to be known beforehand.
@@ 1.3 Estimating the Number of Communities in SBM
Spectral methods tend to rely on few extreme eigenvalues of a suitably modified adjacency matrix to estimate the number of blocks. Computing few extreme eigenvalues is computationally cheap even for large matrices making the spectral methods computationally efficient. One of the popular methods Lei (2016) [30]’s Goodness-of-Fit (GoF) approaches uses the maximum and minimum eigenvalues of the generalized Wigner matrix to estimate a SBM network. Unlike majority of the existing methods, Lei (2016) [30]’s GoF approach estimates the number of communities with $K$ increasing at the rate of $O(n^{1/6-\tau})$ where $\tau>0$. Unfortunately, GoF method is only applicable for dense SBM networks where the degree $d=O(n)$. Other popular spectral approaches are Le and Levina (2022) [31]’s BHMC and NB methods that utilize the spectrum of Bethe-Hessian matrix and Non-backtracking matrix respectively. Empirically, their methods are applicable for both sparse and dense SBMs/DCSBMs but theoretically they were only able to show the validity of their methods for a narrow region. They do allow for the number of communities to increase with $n$ but they do not give an analytical rate. Recently, Chen et al. (2023) [32]’s cross-validated eigenvalues use the idea that the sample eigenvectors under the null hypothesis should be orthogonal to the true latent dimensions. They adapt this idea to consistently estimate $K$ when the degree of the network $d\geq O(log(n))$. Their result holds for fixed $K$ where the $K$ largest eigenvalues satisfy some conditions related to the signal strength.
@@
Organization of the Paper. The rest of the paper is organized as follows. Section 2 introduces the basic set up by introducing the notations used in the paper while borrowing and defining key ideas in the paper. Particularly, we borrow a result on the limiting distribution of the second eigenvalue of appropriately scaled adjacency matrix. This section also introduces the SMT algorithm. Section 3 establishes the theoretical results involving SMT algorithm such as consistency. Section LABEL:applied introduces a new algorithm to extract SBM resembling networks out of scRNA-seq datasets.
@@ 2.2 Main Idea
$M^{(i)}(u,v)=\begin{cases}\frac{1}{\sqrt{N_{i}p_{ii}(1-p_{ii})}},&\text{ if }A^{(i)}_{\star}(u,v)=1,u,v=1,\cdots,N_{i}\\ 0,&\text{ if }A^{(i)}_{\star}(u,v)=0,u,v=1,\cdots,N_{i},\end{cases}$ (2)
@@ 2.3 Second Eigenvalues of Erdös Rényi graphs
Notice that the first eigenvalue of an Erdös Rényi graph goes to infinity as the size of the Erdös Rényi graph increases. We can think of the first eigenvalue as a measure of signal. However, the second eigenvalue of the Erdös Rényi network when appropriately centered and scaled tend to remain stable which can be represented as the noise. We use this observation about the second eigenvalue of Erdös Rényi graph to identify an Erdös Rényi graph. To this end, we collect a pertinent result.
@@ 3.1 Asymptotic Null
[…]
for $K_{\star}^{(n)}=O(log(n))$.
  • Proof
@@
Theorem 3.1 states that the appropriately centered and scaled eigenvalue of the estimated scaled adjacency matrices corresponding to any $K_{\star}^{(n)}$ Erdös Rényi graph converges to the Tracy-Widom distribution. The proof of the above theorem follows from bounding the estimation error using Weyl’s inequality. Note that the above theorem puts an upper bound on $K_{\star}^{(n)}$ for the semi-sparse graph as $O(n)$. The above condition on $K_{\star}^{(n)}$ in (11) gets relaxed for the dense graphs which gives us the following corollary.
@@
For dense graphs when $\hat{G}_{\rho_{n}}(i,i)=O(1)$ with the growing number of communities $K_{\star}^{(n)}=O(n^{1/6-\tau})$ with $\tau>0$, we show that the second largest eigenvalue of the Erdös Rényi converges to the Tracy-Widom distribution under the following conditions
[…]
@@ 3.2.1 Investigations into Type I error and Power
(Consistency of $\hat{K}$) The estimate given for $K_{\star}^{(n)}$ in algorithm 1 converges to $K_{\star}^{(n)}$ provided that the underlying SBM satisfies assumption A1 and one of the following condition is satisfied: i) $K_{\star}^{(n)}=O(n^{1/6-\tau})$ and $G_{\rho_{n}}(i,i)=O(1)$, $\forall i=1,\cdots,K_{\star}^{(n)}$, ii) $K_{\star}^{(n)}=O(log(n))$, $G_{\rho_{n}}(i,i)=O((n/K_{\star}^{(n)})^{-2/3})$ then $\hat{K}$ given in (12) is consistent
@@ 4 Numerical Evaluation
Table 3: Numerical Accuracy Comparison. The following table compiles the numerical accuracy of BHMC, LRBIC, NCV, ECV, StGoF, and SMT for various scenarios with the true number of communities fixed at $2$ and the network size at $500$.
@@
Table 4: Numerical Accuracy Comparison. The following table compiles the numerical accuracy of BHMC, LRBIC, NCV, ECV, StGoF, and SMT for various scenarios with the true number of communities fixed at $3$ and the network size at $500$.
@@
Table 5: Numerical Accuracy Comparison. The following table compiles the numerical accuracy of BHMC, LRBIC, NCV, ECV, StGoF, and SMT for various scenarios with the true number of communities fixed at $4$ and the network size at $500$.
@@
5 SBM networks from scRNA-seq datasets
For applying SMT on a network generated out of scRNA-seq data, we have to carefully design a method to construct SBMs out of scRNA-seq networks. However, most of the data generating process frequently require some fine tuning of free hyperparameters, see Kiselev et al. (2017) [4] among others.
To this end, we perform data filtering and data normalization according to the best practices, see Heumos et al. (2023) [2]. In general, we filter out cells that have less cell counts or more cell counts. Similarly, we filter out ubiquitous genes and/or rare genes. For this kind of filtering, we vary the hyperparameter $\eta$ over a grid of values. Subsequently, we use the correlation matrix to construct a similarity matrix. Then, we construct an adjacency matrix by thresholding using the $\kappa^{th}$ quantile of the correlation matrix $R$,i.e.,
@@ 5 SBM networks from scRNA-seq datasets
For each adjacency matrix, we use SMT to estimate the true number of communities with $\alpha$ varying over $\{0.01,0.05\}$. For convenience, we denote the estimated number of communities as $\hat{K}_{\alpha,\kappa,\eta}$. Then for every adjacency matrix we obtain the complement of the adjacency matrix. Subsequently for every combination of $\{\alpha,\kappa,\eta\}$, we compute the log likelihood under the SBM for each estimate $\hat{K}_{\alpha,\kappa,\eta}$. Finally, we select the network for which the log likelihood is maximized. This selected network would be the network that would closely resemble to a SBM where SMT can be applied. It is worth noting that practitioners may still chose other methods such as BHMC or EigCV instead of SMT. However, SMT has the comparative advantage that it is less sensitive to the out-in ratio making it a suitable candidate for the network analysis of scRNA-seq datasets.
6 Real Data Analysis
First we benchmark the performance of SMT against the competing methods on the six benchmark scRNA-seq datasets where the reference clusters are known. Second, we use our method for subgroup analysis of bipolar scRNA-seq datasets.

### v2

@@ Abstract
Clustering of single-cell RNA sequencing (scRNA-seq) datasets can give key insights into the biological functions of cells. Therefore, it is not surprising that network-based community detection methods are increasingly being used for the clustering of scRNA-seq datasets. The main challenge in implementing network-based community detection methods for scRNA-seq datasets is that these methods apriori require the true number of communities or blocks for estimating the community memberships. Although there are existing methods for estimating the number of communities, they are not robust for noisy scRNA-seq datasets. Moreover, we require an appropriate method for extracting suitable networks from scRNA-seq datasets. For addressing these issues, we present a two-fold solution: i) a simple likelihood-based approach for extracting stochastic block models (SBMs) out of scRNA-seq datasets, ii) a new sequential multiple testing (SMT) method for estimating the number of communities in SBMs. Additionally, the main challenge with existing data analyses of scRNA-seq datasets is that they are sensitive to the choice of hyper-parameters. Our proposed approach also adresses this.
We study the theoretical properties of SMT and establish its consistency under moderate sparsity conditions. We compare the numerical performance of the SMT with several existing methods. We also show that our approach performs competitively well against existing methods for estimating the number of communities on benchmark scRNA-seq datasets. Finally, we use our approach for estimating subgroups of a human retina bipolar single cell dataset. In the above instances, we never fix our hyper-parameters but rather select the best set of hyper-parameters through a grid-search.
@@ 1.1 Motivation
Recently, several authors have proposed multiple approaches to tackle the problem of clustering of noisy scRNA-seq datasets. Broadly these ideas can be organized into the following steps: i) filtering low-quality cells and genes, ii) log-normalizing the read counts, iii) reducing the number of features or performing dimension reduction, iv) estimating the number of communities, and v) clustering the transformed data; see Heumos et al. (2023) [18] for a thorough discussion. The first step of filtering low-quality genes and cells are required for reducing the noise in scRNA-seq datasets whereas log-normalizing in the second step is required for boosting the signal strength while facilitating an even comparison of cells. Unfortunately, the first two steps cannot completely reduce the elevated noise level from scRNA-seq datasets and further steps are warranted. The existing methodology deals with the remaining noisiness of scRNA-seq datasets either through dimension reduction methods or by estimating the number of blocks. However, the existing approaches require fine-tuning of hyperparameters, see Shekhar et al. (2016) [42], Kiselev et al. (2017) [23]. For instance, Louvain and Leiden algorithms implemented in Seurat (Satija et al. (2015) [39]), require the number of nearest neighbors beforehand to construct a weighted network for which they estimate the number of communities by modularity maximization approach (Newman and Girvan (2004) [35]). These approaches require user-informed fine tuning of hyperparameters result in a clustering mechanism that is sensitive to hyperparameters.
A clustering process that divides the cells into multiple groups is extremely desirable as it helps us break-down the cell network into multiple homogeneous groups of similar cellular functions. This could give us insight into the latent biological process. This goal goes in hand with the the goal of developing clustering process that does not require frequent fine-tuning of hyperparameters and are fast algorithms. With these goals in mind, we study stochastic block models for scRNA-seq datasets.
@@ 1.2 Stochastic Block Models
The stochastic block model (SBM) proposed by Holland et al. (1983) [19] is one of the popular models for network data with the block or community structure. SBM simplifies the network model by assuming that the edge probability between any pair of nodes is solely determined by their respective community memberships. Although SBM has found favor in many application areas, it is still somewhat not adequate in the other application domains. Recently many extensions and variants of SBM were proposed. For instance, Airoldi et al. (2008) [1] proposed the mixed membership stochastic block model allowing nodes to belong to multiple networks. Karrer and Newman (2011) [22] proposed degree corrected stochastic block models (DCSBMs) relaxing the assumption of homogeneity of nodes. Sischika and Kauermann (2025) [43] combined a stochastic block model with a Graphon model. Currently, there are many existing methods for estimating the community structure of SBMs including modularity maximization (Newman and Girvan (2004) [35]), Louvain modularity algorithms (Blondel et al. (2008) [6]), likelihood-based approaches (Bickel and Chen (2009) [5], Zhao et al. (2012) [47], Choi et al. (2012) [13], Amini et al. (2013) [3]), spectral clustering methods (Rohe et al. (2011) [38], Lei and Ronaldo (2015) [29], Joseph and Yu (2016) [21], Sengupta and Chen (2015) [41]) amongst others, see Zhao et al. (2017) [48] for a review. Several existing methods such as Newman and Girvan (2004) [35], Bickel and Chen (2009) [5], Zhao et al. (2012) [47], Qin and Rohe (2013) [36], Amini et al. (2013) [3] have also shown to be consistent for both sparse and dense DCSBMs. However, invariably all of the existing community detection methods for SBMs require the true number of communities to be known beforehand.
@@ 1.3 Estimating the Number of Communities in SBM
Spectral methods tend to rely on few extreme eigenvalues of a suitably modified adjacency matrix to estimate the number of blocks. Computing few extreme eigenvalues is computationally cheap even for large matrices making the spectral methods computationally efficient. One of the popular methods Lei (2016) [30]’s Goodness-of-Fit (GoF) approaches uses the maximum and minimum eigenvalues of the generalized Wigner matrix to estimate a SBM network. Unlike majority of the existing methods, Lei (2016) [30]’s GoF approach estimates the number of communities with $K$ increasing at the rate of $O(n^{1/6-\tau})$ where $\tau>0$. Unfortunately, GoF method is only applicable for dense SBM networks where the degree $d=O(n)$. Other popular spectral approaches are Le and Levina (2022) [26]’s BHMC and NB methods that utilize the spectrum of Bethe-Hessian matrix and Non-backtracking matrix respectively. Empirically, their methods are applicable for both sparse and dense SBMs/DCSBMs but theoretically they were only able to show the validity of their methods for a narrow region. They do allow for the number of communities to increase with $n$ but they do not give an analytical rate. Recently, Shao and Le [49] proposed a non-backtracking matrix based approach for estimating the number of communities in sparse and imbalanced settings. However, they provided empirical and theoretical justification for only the Erdös Rényi case. Also, Chen et al. (2023) [11]’s cross-validated eigenvalues use the idea that the sample eigenvectors under the null hypothesis should be orthogonal to the true latent dimensions. They adapt this idea to consistently estimate $K$ when the degree of the network $d\geq O(log(n))$. Their result holds for fixed $K$ where the $K$ largest eigenvalues satisfy some conditions related to the signal strength.
@@
Organization of the Paper. The rest of the paper is organized as follows. Section 2 introduces the basic set up by introducing the notations used in the paper while borrowing and defining key ideas in the paper. Particularly, we borrow a result on the limiting distribution of the second eigenvalue of appropriately scaled adjacency matrix. This section also introduces the SMT algorithm. Section 3 establishes the theoretical results involving SMT algorithm such as asymptotic null, asymptotic power, and consistency. Section 4 performs simulations and compares our method against other existing methods. Section 5 introduces a new algorithm to extract SBM resembling networks out of scRNA-seq datasets. The section also performs the benchmark data analysis and real data analysis. Section 6 summarizes the paper.
@@ 2.2 Main Idea
[…]
@@ 2.3 Second Eigenvalues of Erdös Rényi graphs
Notice that the first eigenvalue of an Erdös Rényi graph goes to infinity as the size of the Erdös Rényi graph increases. We can think of the first eigenvalue as a measure of signal. However, the second eigenvalue of the Erdös Rényi network when appropriately centered and scaled tend to remain stable which can be represented as the noise. We use this observation about the second eigenvalue of Erdös Rényi graph to identify an Erdös Rényi graph. To this end, we collect a pertinent result from [27].
@@ 3.1 Asymptotic Null
$(n/K_{\star}^{(n)})^{2/3}(\lambda_{2}(\hat{M}^{(i)})-2-\frac{1}{\hat{\mu}_{i}})\xrightarrow[]{D}TW_{1}(\cdot),$ (10)
when $\hat{G}_{\rho_{n}}(i,i)\geq(n/K_{\star}^{(n)}))^{-2/3},\forall i=1,\cdots,K_{\star}^{(n)}$ for $K_{\star}^{(n)}=O(log(n))$.
@@
Theorem 3.1 states that the appropriately centered and scaled eigenvalue of the estimated scaled adjacency matrices corresponding to any $K_{\star}^{(n)}$ Erdös Rényi graph converges to the Tracy-Widom distribution. The proof of the above theorem follows from bounding the estimation error using Weyl’s inequality. Note that the above theorem puts an upper bound on $K_{\star}^{(n)}$ for the semi-sparse graph as $O(n)$. The above condition on $K_{\star}^{(n)}$ in (11) gets relaxed for the dense graphs to the $O(n^{1/4-\tau})$.
@@
For dense graphs when $\hat{G}_{\rho_{n}}(i,i)=O(1)$ with the growing number of communities $K_{\star}^{(n)}=O(n^{1/4-\tau})$ with $\tau>0$, we show that the second largest eigenvalue of the Erdös Rényi converges to the Tracy-Widom distribution under the following conditions
$(n/K_{\star}^{(n)})^{(2/3)}(\lambda_{2}(\hat{M}^{(i)})-2-\frac{1}{\hat{\mu}_{i}})\xrightarrow[]{D}TW_{1}(\cdot),$ (11)
when $\hat{G}_{\rho_{n}}(i,i)=O(1),\forall i=1,\cdots,K_{\star}^{(n)}$.
@@ 3.2.1 Investigations into Type I error and Power
(Consistency of $\hat{K}$) The estimate given for $K_{\star}^{(n)}$ in algorithm 1 converges to $K_{\star}^{(n)}$ provided that the underlying SBM satisfies assumption A1 and one of the following condition is satisfied: i) $K_{\star}^{(n)}=O(n^{1/4-\tau})$ and $G_{\rho_{n}}(i,i)=O(1)$, $\forall i=1,\cdots,K_{\star}^{(n)}$, $i$) $K_{\star}=O(log(n)),G_{\rho_{n}}(i,i)=O((n/K_{\star}^{(n)})^{-2/3})$ then $\hat{K}$ given in (12) is consistent
@@ 4 Numerical Evaluation
Table 3: Numerical Accuracy Comparison. The following table compiles the numerical accuracy of BHMC, LRBIC, NCV, ECV, StGoF, and SMT for various scenarios with the true number of communities fixed at 22 and the network size at $500$.
@@
Table 4: Numerical Accuracy Comparison. The following table compiles the numerical accuracy of BHMC, LRBIC, NCV, ECV, StGoF, and SMT for various scenarios with the true number of communities fixed at 33 and the network size at $500$.
@@
Table 5: Numerical Accuracy Comparison. The following table compiles the numerical accuracy of BHMC, LRBIC, NCV, ECV, StGoF, and SMT for various scenarios with the true number of communities fixed at 44 and the network size at $500$.
@@
5 Real Data Analysis
We adapt our method for performing clustering of the scRNA-seq datasets. This begs a question whether SBM is the best model for clustering of the scRNA-seq datasets? The case may be easily made that they are not the best model for the clustering of a particular scRNA-seq dataset. However, the SBM has clear advantage in being the simplest, the parsimonious and most explainable model for clustering of scRNA-seq datasets. This has the added advantage that the parsimonious model tend to be more robust and when the data is as noisy as scRNA-seq dataset then there is a case for developing robust and parsimonious solutions. We believe that SBM based clustering method has the explainable power of decomposing cell networks into multiple homogeneous groups which would encourage many practitioners to adopt them. Additionally, we also validate our method by tracking the performance of our methods against existing state of art methods on the benchmark datasets. For the practical implementation, we note that Grabski, Street and Irizarry (2024)[17] remarked that most of the clustering methods of scRNA-seq are heuristic in nature and does not account for statistical uncertainty. We attempt to approach this by removing any fine-tuning of the hyperparameters and developing a model based clustering of the scRNA-seq datasets.
Extracting SBM networks from scRNA-seq datasets. For applying SMT on a network generated out of scRNA-seq data, we have to carefully design a method to construct SBMs out of scRNA-seq networks. However, most of the data generating process frequently require some fine tuning of free hyperparameters, see Kiselev et al. (2017) [23], Grabski et al. (2024) [17] among others. To this end, we perform data filtering and data normalization according to the best practices, see Heumos et al. (2023) [18]. In general, we filter out cells that have less cell counts or more cell counts. Similarly, we filter out ubiquitous genes and/or rare genes. In particular, we vary the hyperparameter $\eta$ over a grid of values. Subsequently, we use the correlation matrix to construct a similarity matrix. Then, we construct an adjacency matrix by thresholding using the $\kappa^{th}$ quantile of the correlation matrix $R$,i.e.,
@@ 5 Real Data Analysis
For each adjacency matrix, we use SMT to estimate the true number of communities with $\alpha$ varying over $\{0.01,0.05\}$. For convenience, we denote the estimated number of communities as $\hat{K}_{\alpha,\kappa,\eta}$. Then for every adjacency matrix we obtain the complement of the adjacency matrix. Subsequently for every combination of $\{\alpha,\kappa,\eta\}$, we compute the log likelihood under the SBM for each estimate $\hat{K}_{\alpha,\kappa,\eta}$. Finally, we select the network for which the log likelihood is maximized. This selected network would be the network that would closely resemble to a SBM. The above approach is a standard approach for any model-based clustering. It is worth noting that practitioners may still chose other methods such as BHMC or EigCV instead of SMT. However, SMT has the comparative advantage that it is less sensitive to the out-in ratio making it a suitable candidate for the network analysis of scRNA-seq datasets.